In a triangle,AD is a 2m high chalkboard which is 1 m above eye level BC,of a learner at C. BC=y.the angle of the top of the chalkboard,from C is B. ACD=aProve that y=2cosB.cos(B-a)/sina

Grade:12

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Kindly upload the question with proper notation or an image of the question so that it becomes clear to understand and solve

Regards
Sripad Sambrani
4 years ago
Given: AD=2m, AB=1m, BC=y, ∟ACB=b, ∟ACD=a
 
RTP: y=2*cosb*cos(b-a)/sina
 
Solution:
BD=(AD-AB)=2-1=1m, BD=AB=1m.
∟ACB=∟DCB, as AB=BD. hence ∟a=2∟b.
In triangle ABC, BC/AB=cotb
=> y/1=cotb
> y=cotb
 
RHS = 2*cosb*cos(b-a)/sina
= 2*cosb*cos(-b)/sin2b [since a=2b, b-2b=-b]
= 2*cosb*cosb/(2*sinb*cosb) [since cos(-θ)=cosθ]
= cotb = y = LHS Thus proved.
 
Trust this helps,
Regards,
SripadS